.2. The historical data of a company’s annual sales is normally distributed with its mean, $330 million and its standard deviation, $45 million. Find the probability that the company’s sales will be between $300 million and $ 400 next year?
Answer #1

Given

= 330

= 45

Now calculate P(300 < x < 400)e

We know that

Z = (x - ​​​​​​) /

P(300 < x < 400)

= P(300-330/45 < z < 400-330/45)

= P(-0.67 < z < 1.56)

= P(0 < z < 0.67) + P(0 < z < 1.56)

= 0.2486 + 0.4406

= 0.6892

P( 300 < x < 400) = 0.6892

The probability that company's sales will be between $300 and $400 is 0.6892.